# Incompatibility of rank-one POVMs and uncertainty relations

### v1

@@ 5. Ghobber-Jaming uncertainty relation for frames
Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ be two frames in $\mathscr{H}$, with lower and upper frame bounds $\alpha_{1},\beta_{1}$ and $\alpha_{2},\beta_{2}$, respectively. Suppose that for any subsets $S\subseteq\{1,\dots,m\}$ and $T\subseteq\{1,\dots,n\}$, satisfying $S\subseteq T$, the condition
$\frac{\mu^{2}}{\alpha_{1}}<\frac{\beta_{2}}{T}$
holds, where $\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|$. Then for all $x\in H$,
$C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}.$
@@
$x_{1}=\sum_{k\in S}\langle x,a_{k}\rangle a_{k}^{*},\quad x_{2}=\sum_{k\in S^{c}}\langle x,a_{k}\rangle a_{k}^{*},$
in which $a_{k}^{*}$ is the canonical dual frame.
$\begin{split}\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j\in T}\left|\sum_{k\in S}\langle x,a_{k}\rangle\langle a_{k}^{*},b_{j}\rangle\right|^{2}\\ \leq\sum_{j\in T}\left(\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}\right)\left(\sum_{k\in S}\left|\langle a_{k}^{*},b_{j}\rangle\right|\right)\\ \leq\mu^{2}\alpha_{2}|T|\|x\|^{2}.\end{split}$
Thus,
$\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j=1}^{n}\left|\langle x_{1},b_{j}\rangle\right|^{2}-\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}\geq\beta_{1}\|x_{1}\|^{2}-\alpha_{2}\mu^{2}\mu^{2}|T|\|x\|^{2}$
This implies
$\begin{split}\beta_{1}^{\frac{1}{2}}\|x_{1}\|&\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}+\alpha_{2}\mu^{2}|T|\|x\|^{2}\right)^{\frac{1}{2}}\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x_{2},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\beta_{2}^{\frac{1}{2}}\|x_{2}\|+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|.\end{split}$
@@
$\begin{split}\alpha_{1}\|x_{2}\|^{2}&\leq\sum_{k=1}^{m}|\langle x_{2},a_{k}\rangle|^{2}\leq\sum_{k=1}^{m}\left(\sum_{i\in S^{c}}|\langle x,a_{i}\rangle|^{2}\right)\left(\sum_{i\in S^{c}}|\langle a_{i}^{*},a_{k}\rangle|^{2}\right)\\ \leq m^{2}M(A^{*},A)^{2}\sum_{k\in s^{c}}|\langle x,a_{k}\rangle|^{2}\end{split}$
Thus we have
$\begin{split}\|x\|\leq\|x_{1}\|+\|x_{2}\|&\leq\frac{1}{\beta_{1}^{\frac{1}{2}}}\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{1/2}\right)\|x_{2}\|+\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\|x\|\end{split}$
we obtain
$\begin{split}C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{1/2}\end{split}$
@@
$C(S,T)=\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\right)\left(\max\{\frac{1}{\beta_{1}^{1/2}},\left\{\beta_{1}^{-\frac{1}{2}},\left((\frac{\beta_{2}}{\alpha_{1}\beta_{1}})^{\frac{1}{2}}+\alpha_{1}^{-\frac{1}{2}}\right)mM(A^{*},A)\right\}\right)^{-1}.$
6. Incompatibility of multiple bases
We consider the incompatibility of multiple bases in this section.
@@ 6. Incompatibility of multiple bases
Orthonormal bases $A_{k}=\{\ket{a^{(k)}_{j}}\}_{j=1}^{n}$ are $s$-order incompatible if:
@@
    For any nonempty $S_{k}\subset A_{k}$ with $\sum_{k=1}^{n}|S_{k}|<s$, $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}=\{0\}.$
@@
    There exist subsets $S_{k}\subset A_{k}$, such that $\sum_{k=1}^{n}|S_{k}|=s$ and $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}\neq\{0\}.$
@@
For $s$-order incompatible bases $\{A_{k}\}_{k=1}^{n}$, any non-zero pure state |$\ket{\varphi}$ if and only if
@@
Let
$n_{\min}=\min_{\ket{\varphi}\neq 0}\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi}).$
Then if $n_{\min}=s$, there exists a pure state |$\ket{\psi}$ such that
$\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi})=s,$
and there does not exist a pure state |$\ket{\psi^{\prime}}$ such that $\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi^{\prime}})<s$.
For such |$\ket{\psi}$, there exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi})$, and |$\ket{\psi}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. The nonexistence of such |$\ket{\psi^{\prime}}$ implies that there do not exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi^{\prime}})$, and |$\ket{\psi^{\prime}}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. These two conditions precisely coincide with (1.1) and (1.2) in Definition 1. Then the claim follows. $\square$
@@ 7. Conclusion
By equating $s$-order incompatibility with the spark, we unify quantum measurement theory with compressed sensing. Key contributions include:
  • POVM Incompatibility: Generalized to non-projective measurements via tight frames.
  • Sharp Uncertainty Bounds: Derived from the spark, improving coherence-based estimates.
  • Multi-Basis Framework: Extended support uncertainty to multi-observable systems.
Future directions include applications in quantum tomography, cryptographic protocol design, and infinite-dimensional spark generalizations.

### v2

@@ 5. Ghobber-Jaming uncertainty for frames
Let $A$ and $B$ be two frames in $\mathscr{H}$, with lower and upper frame bounds $\alpha_{1},\beta_{1}$ for $A$ and $\alpha_{2},\beta_{2}$ for $B$, respectively. Suppose that for any subsets $S\subseteq\{1,\dots,m\}$ and $T\subseteq\{1,\dots,n\}$, satisfying
$|S||T|<\frac{\beta_{1}}{\alpha_{2}}\cdot\frac{1}{M(A^{*},B)^{2}}.$
Then, for any $x\in\mathscr{H}$, we have
$C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}},$
where
$\begin{split}C(S,T)=&\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\right)\\ &\times\left(\max\left\{\beta_{1}^{-1/2},\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{\frac{1}{2}}\right)\alpha_{1}^{-\frac{1}{2}}mM(A^{*},A)\right\}\right)^{-1}.\end{split}$
@@
$x_{1}=\sum_{k\in S}\langle x,a_{k}\rangle a_{k}^{*},\quad{\rm and}\quad x_{2}=\sum_{k\in S^{c}}\langle x,a_{k}\rangle a_{k}^{*},$
in which $a_{k}^{*}$ is the canonical dual frame. For $x_{1}$, we have
$\begin{split}\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}&=\sum_{j\in T}\left|\sum_{k\in S}\langle x,a_{k}\rangle\langle a_{k}^{*},b_{j}\rangle\right|^{2}\\ &\leq\sum_{j\in T}\left(\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}\right)\left(\sum_{k\in S}\left|\langle a_{k}^{*},b_{j}\rangle\right|^{2}\right)\\ &\leq\alpha_{2}M(A^{*},B)^{2}|S||T|\|x\|^{2}.\end{split}$
Therefore,
$\begin{split}\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}&=\sum_{j=1}^{n}\left|\langle x_{1},b_{j}\rangle\right|^{2}-\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}\\ &\geq\beta_{1}\|x_{1}\|^{2}-\alpha_{2}M(A^{*},B)^{2}|S||T|\|x\|^{2}.\end{split}$
It yields that
(5.1) $\begin{split}\beta_{1}^{\frac{1}{2}}\|x_{1}\|\leq&\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}+\alpha_{2}M(A^{*},B)^{2}|S||T|\|x\|^{2}\right)^{\frac{1}{2}}\\ \leq&\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|\\ \leq&\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x_{2},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}\\ &+\alpha_{2}^{\frac{1}{2}}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|\\ \leq&\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\beta_{2}^{\frac{1}{2}}\|x_{2}\|+\alpha_{2}^{\frac{1}{2}}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|.\end{split}$
@@
(5.2) $\begin{split}\alpha_{1}\|x_{2}\|^{2}&\leq\sum_{k=1}^{m}|\langle x_{2},a_{k}\rangle|^{2}\leq\sum_{k=1}^{m}\left(\sum_{i\in S^{c}}|\langle x,a_{i}\rangle|^{2}\right)\left(\sum_{i\in S^{c}}|\langle a_{i}^{*},a_{k}\rangle|^{2}\right)\\ &\leq m^{2}M(A^{*},A)^{2}\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}.\end{split}$
By (5.1) and (5.2), we have
$\begin{split}\|x\|\leq\|x_{1}\|+\|x_{2}\|\leq&\frac{1}{\beta_{1}^{\frac{1}{2}}}\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{1/2}\right)\|x_{2}\|\\ &+\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|.\end{split}$
Regrouping terms, we obtain
$\begin{split}C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{1/2},\end{split}$
@@
$\begin{split}C(S,T)=&\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\right)\\ &\times\left(\max\left\{\beta_{1}^{-1/2},\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{\frac{1}{2}}\right)\alpha_{1}^{-\frac{1}{2}}mM(A^{*},A)\right\}\right)^{-1}.\end{split}$
Remark 5.2.
It is interesting to further improve the bound in terms of
$\mu_{r}(A,B)=\max_{1\leq j\leq n}\left(\sum_{k=1}^{m}|\langle a_{k},b_{j}\rangle|^{r^{\prime}}\right)^{\frac{r}{r^{\prime}}},$
where $r\in[1,2]$ and $11/r+1/r^{\prime}=$ 1, similarly as done in [15].
6. Incompatibility of multiple tight frames
The main goal of this section is to extend the notion of $s$-order incompatibility for multiple tight frames ($n\geq 3$). When dealing with multiple frames, a suitable approach should consider the overall coherence of the entire collection, rather than analyzing each pair separately and subsequently combining the results. To illustrate the differences, we conclude this section with an example.
@@ 6. Incompatibility of multiple tight frames
Suppose $A_{i}=\{a_{k}^{(i)},1\leq k\leq m_{i}\}$ are $n$ tight frames. They are said to be $s$-order incompatible if:
@@
    For any non-empty subsets $S_{i}\subset I_{i}$ satisfying $\sum_{i=1}^{n}|S_{i}|<s$, and for any non-zero $x\in\mathscr{H}$, the following
    $\sum_{k\in S_{i}}\left|\left\langle x,a_{k}^{(i)}\right\rangle\right|^{2}=\alpha_{i}\|x\|^{2},$
    cannot hold for all $i=1,2,\ldots,n$.
@@
    There exist non-empty subsets $S_{i}\subseteq I_{i}$ with $\sum_{i=1}^{n}|S_{i}|=s$, and a non-zero $x\in\mathscr{H}$, such that
    $\sum_{k\in S_{i}}\left|\left\langle x,a_{k}^{(i)}\right\rangle\right|^{2}=\alpha_{i}\|x\|^{2}$
    holds for all $i=1,2,\ldots,n$.
@@
The tight frames $A_{1},A_{2},\cdots,A_{n}$ are $s$-order incompatible if and only if for any non-zero pure state |$\ket{\varphi}$, it holds that
@@
Denote
$n_{\min}=\min_{\ket{\varphi}\neq 0}\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi})=t.$
Then there exists a pure state |$\ket{\psi}$ such that
$\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi})=t,$
and there does not exist a pure state |$\ket{\psi^{\prime}}$ such that $\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi^{\prime}})<t$.
For such |$\ket{\psi}$, there exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi})$, and |$\ket{\psi}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. The nonexistence of such |$\ket{\psi^{\prime}}$ implies that there do not exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi^{\prime}})$, and |$\ket{\psi^{\prime}}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. These two conditions precisely coincide with the two conditions in Definition 6.1, which implies that $t=s$. Then the claim follows. $\square$
We show that the multiple cases ($n\geq 3$) can not be derived from the classical two-frame case.
Theorem 6.3.
Suppose tight frames $A_{1},A_{2},\cdots,A_{n}$ are $s$-order incompatible and each pair $A_{i}$ and $A_{j}$ are $s_{ij}$-order incompatible, then we have
(6.1) $\frac{1}{2}\sum_{i<j}s_{ij}\leq s.$
Proof.
Since $A_{1},A_{2},\cdots,A_{n}$ are $s$-order incompatible, then there exists a nonzero $x\in\mathscr{H}$ such that
(6.2) $n_{A_{1}}(x)+n_{A_{2}}(x)+\ldots+n_{A_{n}}(x)\geq s.$
On the other side, since $A_{i}$ and $A_{j}$ are $s_{ij}$-order incompatible, we have
(6.3) $n_{A_{i}}(x)+n_{A_{j}}(x)\geq s_{ij}.$
Combining (6.2) and (6.3), we get the result.
The following example shows that the strict inequality in (6.1) can occur.
Example 6.4.
Consider the frames $A=\{a_{1},a_{2},a_{3}\}$, $B=\{b_{1},b_{2},\ldots,b_{5}\}$ and $C=\{c_{1},c_{2},c_{3},c_{4}\}$, where
  1. (1)
    $a_{1}=(2\sqrt{3},1),a_{2}=(-2,-\sqrt{3}),a_{3}=(0,2\sqrt{3});$
  2. (2)
    $b_{1}=(1,1),b_{2}=(\frac{1}{2},\frac{1}{2}),b_{3}=(0,1),b_{4}=(2,-\sqrt{3}),$
    $b_{5}=(-\left(2\sqrt{3}-\frac{5}{4}\right)^{\frac{1}{2}},-\left(2\sqrt{3}-\frac{5}{4})^{\frac{1}{2}}\right)$;
  3. (3)
    $c_{1}=(2,-\sqrt{3}),c_{2}=(2\cdot 3^{\frac{1}{4}},2\cdot 3^{\frac{1}{4}}),c_{3}=(-2,\sqrt{3}),c_{4}=(0,-\sqrt{2})$.
(a) (1)
(b) (2)
(c) (3)
Figure 2. frames vectors of $A,B$ and $C$
Straightforward computations show that $s_{AB}=5$, $s_{BC}=5$, $s_{AC}=4$ and $s=8$. Therefore,
$\frac{s_{AB}+s_{BC}+s_{AC}}{2}<s.$
This means that the cases for multiple measurements $n\geq 3$ differs from the uncertainty relation for two bases.
@@ 7. Conclusion
In this paper, we show that the $s$-order incompatibility of two orthonormal bases coincides with the spark of matrix formed by the measurement vectors, thereby linking quantum measurement with compressed sensing. The main contribution of this work is the classification of the incompatibility of two and multiple ($n\geq 3$) tight frames (rank-one POVMs), leading to sharp support uncertainty relations and new generalizations of Ghobber-Jaming uncertainty relation. Potential directions for future research include applications in quantum tomography, cryptographic protocol design, and explicit calculations of the order of incompatibility.
Acknowledgments
This work was partially supported by NSFC Grant No.12101451.